Normal levels of cholesterol, C27H46O, in blood serum are around 150mg in 100cm3. Convert this to %m/V and molarity.
1
Expert's answer
2020-07-29T06:39:19-0400

1) Concentration of cholesterol in m/V is:

C(C27H46O) = 150 mg/ 100 cm3 = 1.5 g/l

2) Molar concentration of cholesterol is:

M (C27H46O)=386 g/mol

n (C27H46O)=1.5/386=0.004 mol

CM(C27H46O)=0.004/1 = mol/l